### v2

@@ A.2 Details of Dataset Construction
Fine-grained Filtering. To further ensure the quality of our data, we conducted additional audio and visual analyses. For the audio analysis, raw videos may contain audio segments without music, such as speech, silence, etc. To ensure the final dataset consists of high-quality video-music pairs, we retain only those videos with a larger portion of music content. We utilize the sound event detection model PANNs (Kong et al., 2020), which provides frame-level event labels across the entire video to identify music events. Based on the observation from a subset of videos, we define two thresholds, i.e., a confidence threshold and a duration threshold, for analyzing the music event. The confidence threshold is set at $0.5$, indicating an audio frame is considered a music event if the PANNs model predicts the probability of the “Music” label to be over $0.5$. The duration threshold of a music event requires that at least $50\%$ of the audio’s frames are classified as music events for the video to be considered valid.
For the visual analysis, some videos only consisting of static images will be removed. Specifically, we uniformly sample multiple temporal windows without overlap from the video. Within each window, we use Structural Similarity Index Measure (SSIM) (Wang et al., 2004) between the first frame and the last frame. By aggregating average SSIM values from all temporal windows, we remove the videos with average SSIM values lower than a threshold of 0.8, empirically.
Music Source Separation. Since the irrelevant human speech in videos poses a negative impact on music generation, we apply music source separation to process the videos. We employ Demucs (Défossez et al., 2019) as the music source separation model to filter out the speech signals.
Audio-Video Alignment Ranking. ImageBind-AV (Girdhar et al., 2023) scores usually reflect the semantic correlation between the vision and audio modality. To construct a high-quality subset with better alignment, we compute the ImageBind-AV scores for all the data and rank them accordingly.
@@
A.3 Details of Evaluation Metrics
@@
Frechet Audio Distance (FAD) [^3] https://github.com/haoheliu/audioldm_eval (Kilgour et al., 2018) (Kilgour et al., 2018) is a reference-free evaluation metric for assessing audio quality. Similar to Frechet Image Distance (FID)(Heusel et al., 2017), it compares the embedding statistics of the generated audio clip with ground truth audio clips. A shorter distance usually denotes better human-perceived acoustic-level audio quality. However, this metric cannot reflect semantic-level information in audio. We report the FAD based on the VGGish(Hershey et al., 2017) feature extractor.
@@
Frechet Distance (FD)33footnotemark: 3 measures the similarity between generated samples and target samples in audio generation fields. It’s similar to FAD but uses a PANNs feature extractor instead. PANNs(Kong et al., 2020) have been pre-trained on AudioSet(Gemmeke et al., 2017), one of the largest audio understanding datasets, thus resulting in a more robust metric than FAD.
@@
Kullback-Leibler Divergence (KL)33footnotemark: 3 reflects the acoustic similarity between the generated and reference samples to a certain extent. It is computed over PANNs’ multi-label class predictions.
@@
Density and Coverage [^4] https://github.com/clovaai/generative-evaluation-prdc (Naeem et al., 2020) measures the fidelity and diversity aspects of the generated samples. Fidelity measures how closely the generated samples match the real ones, while diversity assesses whether the generated samples capture the full range of variation found in real samples. We use CLAP(Wu et al., 2023b) embeddings for manifold estimation.
@@
Imagebind Score [^5] https://github.com/facebookresearch/ImageBind (Girdhar et al., 2023) assesses to what extent the generated music aligns with the videos. Despite the fact that Imagebind extends the CLIP to six modalities, we only use the branches of audio and vision. Since we use ImageBind to filter out video-audio pairs with a low matching score during dataset construction, the ImageBind score is naturally used in our evaluation. We acknowledge that ImageBind is not specifically trained on music data, which may limit its effectiveness in capturing the full complexity of video-music alignment. However, it remains the most suitable option available for this task at present.
A.4 Details of the Inference Process
@@
When predicting music on videos of arbitrary length, maintaining music consistency and coherence is particularly important. However, it leads to a significant challenge on computational resources due to the quadratic dependency of transformers-based models on sequence length (Zaheer et al., 2020; Beltagy et al., 2020). To address this problem, we adopt a sliding window approach for inferring the whole video.
@@
During inference, given an input video with a length of $L$, we define $L_{s}$ as the length of the sliding window and $O$ as the overlap between consecutive windows. With the window start position $t$ initially set to 00, the inference involves the following steps compactly while $t+L_{s}\leq L$: (1) using a visual encoder to extract feature representations X and capture long-term dependencies $\textbf{X}_{l}$; (2) collecting embeddings within the window $[t,t+L_{s}]$ to obtain $\textbf{X}_{s}$; (3) predicting the music tokens $\bar{\textbf{Y}}$ for the reduced window $[t,t+L_{s}-O]$ based on $\textbf{X}_{l}$ and $\textbf{X}_{s}$; (4) decoding $\bar{\textbf{Y}}$ to the predicted audio $\bar{\textbf{A}}$ using the audio decoder; (5) move the window forward by setting $t=t+L_{s}-O$, and repeating steps (2) to (5) until the end of the video.
@@
After finishing above steps, we can concatenate all musical segments to form a cohesive piece of music that aligns in duration with the video.

### v3

@@ 9 Details of Dataset Construction
Fine-grained Filtering. To further ensure the quality of our data, we conducted additional audio and visual analyses. For the audio analysis, raw videos may contain audio segments without music, such as speech, silence, etc. To ensure the final dataset consists of high-quality video-music pairs, we retain only those videos with a larger portion of music content. We utilize the sound event detection model PANNs [34], which provides frame-level event labels across the entire video to identify music events. Based on the observation from a subset of videos, we define two thresholds, i.e., a confidence threshold and a duration threshold, for analyzing the music event. The confidence threshold is set at $0.5$, indicating an audio frame is considered a music event if the PANNs model predicts the probability of the “Music” label to be over $0.5$. The duration threshold of a music event requires that at least $50\%$ of the audio’s frames are classified as music events for the video to be considered valid.
For the visual analysis, some videos only consisting of static images will be removed. Specifically, we uniformly sample multiple temporal windows without overlap from the video. Within each window, we use Structural Similarity Index Measure (SSIM) [72] between the first frame and the last frame. By aggregating average SSIM values from all temporal windows, we remove the videos with average SSIM values lower than a threshold of 0.8, empirically.
Music Source Separation. Since the irrelevant human speech in videos poses a negative impact on music generation, we apply music source separation to process the videos. We employ Demucs [59] as the music source separation model to filter out the speech signals.
Audio-Video Alignment Ranking. ImageBind-AV [22] scores usually reflect the semantic correlation between the vision and audio modality. To construct a high-quality subset with better alignment, we compute the ImageBind-AV scores for all the data and rank them accordingly.
@@
10 Additional Dataset Analysis
Music Genre Distribution. To better understand the diversity of our dataset, we analyze the distribution of music genres across all selected video-music pairs. The results are illustrated in Fig. A1. As shown, the dataset covers a wide range of genres, including but not limited to electronic, classical, pop, and rock. The diversity in genres ensures that the dataset provides a comprehensive foundation for the task of video-to-music generation, enabling robust performance across various musical styles.
Instrument Usage Distribution. We also analyze the usage of different instruments within the dataset. The distribution is shown in Fig. A2. The frequent occurrences of synthesizers, pianos, and drums, along with a variety of other instruments, ensure the ability to capture diverse musical elements in the video-to-music generation task.
Mood Information. In addition to genres and instruments, we also explore the mood information present in the music data. A word cloud representation of the mood labels is shown in Fig. A3, where the font size corresponds to the frequency of each mood label. Commonly occurring moods include inspiring, happy, dark, powerful, and sentimental, showcasing the emotional diversity of the dataset. This emotional richness enhances the dataset’s capacity to generate music that aligns closely with the mood conveyed in videos.
All music-related metadata, including genre, instrument, and mood, is annotated using Qwen2-Audio, a state-of-the-art (SOTA) model for music understanding.
11 Details of Evaluation Metrics
@@
Frechet Audio Distance (FAD) is a reference-free evaluation metric for assessing audio quality. Similar to Frechet Image Distance (FID)[28], it compares the embedding statistics of the generated audio clip with ground truth audio clips. A shorter distance usually denotes better human-perceived acoustic-level audio quality. However, this metric cannot reflect semantic-level information in audio. We report the FAD based on the VGGish[27] feature extractor.
@@
Frechet Distance (FD) measures the similarity between generated samples and target samples in audio generation fields. It’s similar to FAD but uses a PANNs feature extractor instead. PANNs[34] have been pre-trained on AudioSet[21], one of the largest audio understanding datasets, thus resulting in a more robust metric than FAD.
@@
Kullback-Leibler Divergence (KL) reflects the acoustic similarity between the generated and reference samples to a certain extent. It is computed over PANNs’ multi-label class predictions.
@@
Density and Coverage [55] measures the fidelity and diversity aspects of the generated samples. Fidelity measures how closely the generated samples match the real ones, while diversity assesses whether the generated samples capture the full range of variation found in real samples. We use CLAP[75] embeddings for manifold estimation.
@@
Imagebind Score [22] assesses to what extent the generated music aligns with the videos. Despite the fact that Imagebind extends the CLIP to six modalities, we only use the branches of audio and vision. Since we use ImageBind to filter out video-audio pairs with a low matching score during dataset construction, the ImageBind score is naturally used in our evaluation. We acknowledge that ImageBind is not specifically trained on music data, which may limit its effectiveness in capturing the full complexity of video-music alignment. However, it remains the most suitable option available for this task at present.
12 Details of the Inference Process
@@
When predicting music on videos of arbitrary length, maintaining music consistency and coherence is particularly important. However, it leads to a significant challenge on computational resources due to the quadratic dependency of transformers-based models on sequence length [82, 5].
@@
To address this problem, we adopt a sliding window approach for inferring the whole video. During inference, given an input video with a length of $L$, we define $L_{s}$ as the length of the sliding window and $O$ as the overlap between consecutive windows. With the window start position $t$ initially set to 00, the inference involves the following steps compactly while $t+L_{s}\leq L$: (1) using a visual encoder to extract feature representations X and capture long-term dependencies $\textbf{X}_{l}$; (2) collecting embeddings within the window $[t,t+L_{s}]$ to obtain $\textbf{X}_{s}$; (3) predicting the music tokens $\bar{\textbf{Y}}$ for the reduced window $[t,t+L_{s}-O]$ based on $\textbf{X}_{l}$ and $\textbf{X}_{s}$; (4) decoding $\bar{\textbf{Y}}$ to the predicted audio $\bar{\textbf{A}}$ using the audio decoder; (5) move the window forward by setting $t=t+L_{s}-O$, and repeating steps (2) to (5) until the end of the video.
@@
After finishing the above steps, we can concatenate all musical segments to form a cohesive piece of music that aligns in duration with the video.
13 Qualitative Analysis
In Fig. A5, our qualitative analysis highlights specific limitations of CMT, Video2Music, and M2UGen. CMT and Video2Music extract visual cues to generate symbolic music, i.e., MIDI notes. However, CMT’s training strategy for symbolic music generation leads to discontinuities, particularly for slowly changing or static frames, where the model fails to predict symbolic music notes, resulting in periods of silence. Additionally, the approach of predicting MIDI notes and then rendering them into audio, as employed by both CMT and Video2Music, lacks high-frequency content, negatively affecting auditory perception. M2UGen utilizes LLMs to fuse multimodal representation and then project LLMs’ embeddings into music via a text-to-music generation model. However, this approach relies on text embeddings as intermediaries, which causes the loss of visual information and restricts the model’s ability to detect nuanced visual variations. As a result, the music generated by this method usually showcases repetitive musical themes and suffers from a lack of diversity, as evidenced in Fig. A5 and the supplementary videos. The last row of Fig. A5 demonstrates that our Long-Short-Term (LST) approach is capable of generating music that is rich in diversity and semantically consistent with the video.
14 User Study Interface
Fig. A4 illustrates the A/B test interface used during the user study. Participants evaluated the videos based on four criteria: Audio Quality, Video-Music Alignment, Musicality, and Overall Assessment. This interface shows participants comparing two videos side-by-side and selecting the better one for each criterion.
15 Supplementary Videos
For additional insights and demonstrations, we kindly refer readers to our supplementary video for a comprehensive showcase of our method’s performance.